Colby, of mass 74 kg, decides to go ziplining down a zipline that is 100 m long on the hypotenuse and declines at an angle of 10°.  If he starts from rest at what speed will he land at the bottom if friction can be ignored?



1
Expert's answer
2021-06-13T07:09:50-0400

"\sin10°=x/100\to x=100\cdot\sin10°=17.4\ (m)"


"mgx=mv^2/2\to v=\sqrt{2gx}=\sqrt{2\cdot9.8\cdot17.4}\approx18.5\ (m/s)" . Answer